OpenStudy (anonymous):

Identify the vertex for the graph of y = -3x^2 + 6x - 4.

OpenStudy (anonymous):

@undeadknight26

OpenStudy (mathmate):

Do you know how to complete squares?

OpenStudy (anonymous):

no

OpenStudy (mathmate):

Then you would have to take it for granted that the x-value of the vertex is \(x=-b/(2a)\) and evaluate y according to the given function. For \(example\), the vertex of y=x^2-3x+2 is +3/(2)=1.5 y=(1.5^2)-3*1.5+2 =-0.25 so the vertex is at (1.5, -0.25)

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

ty

OpenStudy (mathmate):

You're welcome! :)

OpenStudy (mathmate):

@style2814 I hope you realized that the calculations I made was an example, and does not correspond to your question in hand!